Solve each problem (3x-2)/5=(4x+1)/3
Answer by tinaweber17(1)

(3x-2)/5=(4x+1)/3
LCD= 15 so you need to multiply each side by 15
3(3x-2)=5(4x+1)
distribute
9x-6=20x+5
subtract the 5 from both sides
9x-11=20x
subtract 9x from both sides
-11=11x
divide by 11 to get the x by itself
x=-1